# Towards One-step Causal Video Generation via Adversarial Self-Distillation

### v1

@@ 1 Introduction
To alleviate this limitation, distillation has emerged as a promising direction for accelerating diffusion models (Sauer et al., 2024b; Yin et al., 2024b; Lu et al., 2025). By reducing a multi-step model to a few-step counterpart, distillation greatly improves efficiency while preserving generation quality. Nonetheless, existing distillation objective primarily focuses on aligning the prediction distributions of the few-step student with those of the multi-step teacher. When the student performs only 1- or 2-step generation, the discrepancy from the multi-step teacher becomes excessively large, making direct alignment unstable and causing severe quality degradation (Cheng et al., 2025; Yin et al., 2025; Huang et al., 2025). In other words, the fewer the denoising steps, the larger the semantic and statistical gap to be bridged, which explains why extremely few-step distillation is particularly challenging.
Figure 1: The different alignment strategies of DMD and ASD during distillation. $p_{\text{n-gen}}$ represents the generation distribution of the $n$-step distillation model. The points plotted in the figure indicate the multiple modes (or peaks) of the distribution, highlighting its multimodal nature.
@@
  • We propose a unified adversarial self-distillation strategy that aligns prediction distributions across different denoising steps of the student model, significantly improving few-step generation quality.
@@ 2.2 Adversarial Distillation
Further advancements have sought to preserve the generation trajectory itself, not just the final output. SDXL-Lightning (Lin et al., 2024) and LADD (Sauer et al., 2024a) achieve this by aligning the intermediate denoising states of the student model with those of the teacher model. Our work employs a discriminator to align the generated results of an $n$-step denoising process with those of an $(n+1)$-step process. This approach enhances the model’s performance and consistency across various limited-step scenarios.
@@ 2.3 Score Distillation
Existing methods (Zhou et al., 2024; Lu et al., 2025) also utilize score-based models to achieve distribution matching across different noise states during distillation. For example, Dreamfusion (Poole et al., 2023) and ProlificDreamer (Wang et al., 2023) leverage a T2I diffusion model’s score function to guide text-conditioned 3D generation. Distribution Matching Distillation (DMD) (Yin et al., 2024b; a) extends this approach to accelerate diffusion model generation by employing the original model as a score estimator for real data and a fine-tuned version as a score estimator for generated data.
@@ 3 Methodology
To achieve high-quality and flexible causal video generation with minimal computational overhead, our distillation process employs two key strategies. We use Distribution Matching Distillation (DMD) to distill a multi-step generative model into a few-step student model through score-based alignment. Meantime, we introduce a novel form of self-supervision by aligning the student model’s $n$-step denoising with its (n+1n\!+\!1)-step version in the distribution level.
This dual supervision not only enhances the generation quality but also enables the final model to produce consistent, hight-quality output across a variety of few-step settings. The few-step generator, the generator’s score estimator and the discriminator are trained alternately to facilitate this process. The pipeline of the Adversarial Self-Distillation (ASD) process is depicted in Fig. 3. During inference, we further improve efficiency and quality by employing a First-Frame Enhancement (FFE) strategy, which uses a more intensive denoising process for the initial frame while reducing the number of steps for subsequent frames.
@@ 3.1 Preliminary: Distribution Matching Distillation
$\approx-\mathop{\mathbb{E}}\limits_{\bm{z},t,\bm{x}_{t}}[s_{\text{data}}\left(\bm{x}_{t},t\right)-s_{\text{gen}}\left(\bm{x}_{t},t\right)\frac{dG_{\theta}(\bm{z})}{d\theta}],$ (2)
where $s_{\text{data}}$ and $s_{\text{gen}}$ are the score functions that trained of real data and generated data, which point towards the higher density of data for $p_{\text{gen},t}$ and $p_{\text{data},t}$. The noisy sample $\bm{x}_{t}$ is obtained by the diffusion forwarding process $\bm{x}_{t}=\alpha_{t}\hat{\bm{x}}_{0}+\sigma_{t}\epsilon,\ \epsilon\sim\mathcal{N}(0,I)$, where $\hat{\bm{x}_{0}}$ is the generated output of the few-step student model $G_{\theta}$. $\alpha_{t},\sigma_{T}>0$ are defined by the noise schedule. $\bm{z}\sim\mathcal{N}(\bm{0},\bm{I})$ is the Gaussian distribution. The true data score $s_{\text{data}}$ is estimated by the original teacher model, while the generation score $s_{\text{gen}}$ is estimated by a "teaching assistant" (TA) model, which is a fine-tuned version of the teacher. The TA model and the few-step generator are typically trained in an alternating fashion. The TA model is fine-tuned on the standard diffusion denoising loss on the generated data from the student model:
@@
2:Initialize few-step student model $G_{\theta}$ with the origianl model
@@
10:  Add noise $\bm{x}_{t}^{2}=(1-t)\bm{x}^{n}_{0}+t\epsilon,\ \epsilon\sim\mathcal{N}(0,I),t\sim\text{Uniform}(\mathcal{T})$
@@ 3.2 Adversarial Self-Distillation
Our core idea is to align the $n$-step denoising distribution with the (n+1n\!+\!1)-step version. This ensures that the model can maintain consistent, high-quality output regardless of the chosen few-step setting. Similar to adversarial diffusion distillation (Sauer et al., 2024b; Lin et al., 2024), this alignment is performed using a discriminator $D_{\psi}^{n}$, which is tasked with making the outputs of adjacent denoising steps indistinguishable through a relativistic pairing GAN objective (Jolicoeur-Martineau, 2018). The adversarial self-distillation loss is represented as:
@@ 3.3 First-Frame Enhancement
Existing causal video generation methods typically employ a uniform number of denoising steps for each frame. However, in causal generation process, the quality of subsequent frames is critically dependent on the preceding frames. The quality defects in earlier frames are highly likely to be propagated throughout the rest of the sequence. This is particularly true for the first frame, which has no prior context and must synthesize a high-quality initial state from a zero-data starting point. Consequently, the generation of the first frame requires dedicated attention to mitigate accumulated error and ensure high overall video quality. Fig. 4 shows the similarity of predicted $\hat{x}_{0}$ for different video frames across various denoising steps. It reveals that the first frame has low similarity across steps, indicating that each denoising step is crucial. In contrast, subsequent frames exhibit higher similarity, suggesting a greater redundancy that makes them more suitable for few-step prediction.
Based on this observation, we propose a novel Frist-Frame Enhancement (FFE) denoising strategy, as detailed in Alg. 2. The first frame undergoes a more intensive denoising process, requiring a minimum of four steps, while subsequent frames can be generated with a significantly reduced number of steps, such as one or two. This frame-based control allows us to enhance the quality of the generated video in the few-step scenario.
@@ 4.1 Implementations
Our approach employs a causal video distillation framework based on the Self Forcing (Huang et al., 2025) training paradigm. The causal model architecture builds upon the Wan2.1-T2V-1.3B (Wang et al., 2025) backbone. We adopt CausVid’s (Yin et al., 2025) initialization protocol to stabilize early causal training phases via asymmetric distillation from a pre-trained bidirectional teacher model. For training data, we utilize the exact text prompts from Self Forcing for fair comparison. The diffusion process is optimized with a 4-step denoising schedule during training, while using our step-skipping strategy during inference to accelerate generation. Adversarial training objective is incorporated via integration of RpGAN (Jolicoeur-Martineau, 2019) objectives with R1 (Mescheder et al., 2018) and R2 regularization terms following R3GAN (Huang et al., 2024a). And we use the frozen fake score function as the backbone of the discriminator following DMD2 Yin et al. (2024a). To reduce computational cost and the number of parameters, we model discriminators $D_{n}$’s output logits as the n-th dimensional output logit of the final layer, thus different $D_{n}$ share the same backbone and classifier head. The performance is rigorously assessed using VBench (Huang et al., 2024b) for multidimensional evaluation and user preference studies to quantify human-perceived visual quality and semantic alignment.
@@ 4.2 Comparison with Existing Baselines
Table 1: Comparison with relevant baselines. We compare our method with representative open-source video generation models of similar parameter sizes and resolutions. For simplicity, n-step generation with FFE strategy is denoted with $n^{*}$. Self Forcing under 1 & 2 step generation is retrained for fair comparison, denoted as Self Forcing.
@@
Figure 6: User preference study. “(n)” indicates the number of denoising steps for each method. Self Forcing under 1 & 2 step generation is retrained for fair comparison, denoted as Self Forcing.
@@
As illustrated in Fig. 5, our method is capable of generating high-fidelity videos using only approximately 8% and 13% of the denoising steps required by Wan 2.1(Wang et al., 2025) and SkyReels(Chen et al., 2025), respectively, while achieving visually superior results. This leads to a substantial reduction in computational cost. Moreover, under the same number of denoising steps, our approach yields improved detail and visual quality compared to CausVid(Yin et al., 2025) and better video details compared to Self-forcing(Huang et al., 2025). Fig. 6 shows the user study results comparing our model against several important baselines. Our approach is consistently preferred over many-step baselines, including the many-step Wan2.1 that our model is initialized from. Meanwhile, our results are 70% better than CausVid (Yin et al., 2025) and on par with Self Forcing. For the extreme 1-step and 2-step generation setting, our method yields a significantly 96% and 62% preference over Self-Forcing, respectively. This demonstrates the effectiveness of our proposed ASD and FFE strategy in improving extremely few-step video generation.
@@ 4.3 Ablation Study
To better understand the contribution of each component to skip-step generation, we conducted a comprehensive ablation study, as shown in Tab. 2. As shown in the first two rows, using identical training data, our ASD training method consistently improves the quality of both 1-step and 2-step generation during inference. For example, it raises the Total Score by 2.52 and the Semantic Score by 6.84 under one-step generation. Comparison between the first and third rows reveals that the FFE inference strategy leads to substantial gains in overall video quality—notably, a 4.19 increase in Total Score and a 10.64 improvement in Semantic Score for one-step generation, even exceeding the quality of the original two-step generation results (row 1). Moreover, combining both methodology further enhances performance, as demonstrated in the last row.
@@ 7 Ethics Statement
This work focuses on improving the efficiency of diffusion-based video generation through distillation. Our contributions are methodological and technical in nature, and do not involve human subjects, personal or sensitive data, or medical applications. The experiments are conducted on publicly available benchmark datasets (e.g., VBench), which are widely used in prior research and do not contain personally identifiable information. We confirm that this research adheres to the ICLR Code of Ethics and complies with all relevant institutional and legal standards. No conflicts of interest or external sponsorship influencing the work are present.
@@ Appendix A Details of Implementations
where $x^{n+1}_{t}\sim p_{\theta,\text{n+1},t}$, $x^{n}_{t}\sim p_{\theta,\text{n},t}$ are the noisy (n+1)-step generation data and n-step generation data, respectively, $\epsilon$ and $\hat{\epsilon}$ are Gaussian noise sampled from $\mathcal{N}(0,1)$, and $f_{\psi}$ is the critic network (discriminator) of GAN. We use $\lambda=600$, $\sigma=0.05$ for all experiments. The generator, fake score estimator, and discriminator are optimized alternately with a ratio of 1:4:1, following(Yin et al., 2024a; Huang et al., 2025). Note that we discard ASD Loss for the last step denoising. For simplicity, we reserve the index n of the n-th discriminator $D_{n}$ that discriminates the $n$-step generation distribution from that of $(n+1)$-step here.
Appendix B Hyperparameter Analysis

### v2

@@ 1 Introduction
To alleviate this limitation, distillation has emerged as a promising direction for accelerating diffusion models (Sauer et al., 2024b; Yin et al., 2024b; Lu et al., 2025). By reducing a multi-step model to a few-step counterpart, distillation greatly improves efficiency while preserving generation quality. Nonetheless, existing distillation objective primarily focuses on aligning the prediction distributions of the few-step student with those of the multi-step teacher. When the student performs only 1- or 2-step generation, the discrepancy from the multi-step teacher becomes excessively large, making direct alignment unstable and causing severe quality degradation (Cheng et al., 2025; Yin et al., 2025; Huang et al., 2025). In other words, the fewer denoising steps, the larger the semantic and statistical gap to be bridged, which explains why extremely few-step distillation is particularly challenging.
Figure 1: Conceptual illustration of the different alignment strategies of DMD and ASD during distillation. The points plotted in the figure indicate the multiple modes (or peaks) of the real data distribution, highlighting its multimodal nature.
@@
  • We propose an adversarial self-distillation strategy that aligns prediction distributions across different denoising steps of the student, significantly improving few-step generation quality.
@@ 2.2 Adversarial Distillation
Further advancements have sought to preserve the generation trajectory itself, not just the final output. SDXL-Lightning (Lin et al., 2024) and LADD (Sauer et al., 2024a) achieve this by aligning the intermediate denoising states of the student model with those of the teacher model. Zhang et al. (2024) and Lin et al. (2025) fine-tune a pre-trained diffusion video model for AR generation via adversarial training, but they directly align one-step outputs with real data. Our work employs a discriminator to align the generated results of an $n$-step denoising process with those of an $(n+1)$-step process. This approach enhances the model’s performance and consistency across various limited-step scenarios.
@@ 2.3 Score Distillation
Existing methods (Zhou et al., 2024; Lu et al., 2025) utilize score-based models to achieve distribution matching across different noise states during distillation. For example, Dreamfusion (Poole et al., 2023) and ProlificDreamer (Wang et al., 2023) leverage a T2I diffusion model’s score function to guide text-conditioned 3D generation. Distribution Matching Distillation (DMD) (Yin et al., 2024b; a) extends this approach to accelerate diffusion model generation by employing the original model as a score estimator for real data and a fine-tuned version as a score estimator for generated data.
@@ 3 Methodology
To achieve high-quality and flexible causal video generation with minimal computational overhead, our distillation process employs two key strategies. We use Distribution Matching Distillation (DMD) to distill a multi-step generative model into a few-step student model through score-based alignment. Meantime, we introduce a novel form of self-supervision by aligning the student model’s $n$-step denoising with its (n+1n\!+\!1)-step version in the distribution level. This dual supervision not only enhances the generation quality but also enables the final model to produce consistent, hight-quality output across a variety of few-step settings. The few-step generator, the generator’s score estimator and the discriminator are trained alternately to facilitate this process. The pipeline of the Adversarial Self-Distillation (ASD) process is depicted in Fig. 3. During inference, we further improve efficiency and quality by employing a First-Frame Enhancement (FFE) strategy, which uses a more intensive denoising process for the initial frame while reducing the number of steps for subsequent frames.
@@ 3.1 Preliminary: Distribution Matching Distillation
$\approx-\mathop{\mathbb{E}}\limits_{\bm{z},t,\bm{x}_{t}}[\left(s_{\text{data}}\left(\bm{x}_{t},t\right)-s_{\text{gen}}\left(\bm{x}_{t},t\right)\right)\frac{dG_{\theta}(\bm{z})}{d\theta}],$ (2)
where $s_{\text{data}}$ and $s_{\text{gen}}$ are the score functions that trained of real data and generated data, which point towards the higher density of data for $p_{\text{gen},t}$ and $p_{\text{data},t}$. The noisy sample $\bm{x}_{t}$ is obtained by the diffusion forwarding process $\bm{x}_{t}=\alpha_{t}\hat{\bm{x}}_{0}+\sigma_{t}\epsilon,\ \epsilon\sim\mathcal{N}(0,I)$, where $\hat{\bm{x}_{0}}$ is the generated output of the few-step student model $G_{\theta}$. $\alpha_{t},\sigma_{T}>0$ are defined by the noise schedule. $\bm{z}\sim\mathcal{N}(\bm{0},\bm{I})$ is the Gaussian distribution.
The true data score $s_{\text{data}}$ is estimated by the original teacher model, while the generation score $s_{\text{gen}}$ is estimated by a "teaching assistant" (TA) model, which is a fine-tuned version of the teacher. The TA model and the few-step generator are typically trained in an alternating fashion. The TA model is fine-tuned on the standard diffusion denoising loss on the generated data from the student model:
@@
2:Initialize few-step student model $G_{\theta}$ with the original model
@@
10:  Add noise $\bm{x}_{t}^{2}=(1-t)\bm{x}^{n}_{0}+t\epsilon,\ \epsilon\sim\mathcal{N}(0,I)$
@@ 3.2 Adversarial Self-Distillation
Our core idea is to align the $n$-step denoising distribution with the (n+1n\!+\!1)-step version. This ensures that the model can maintain consistent, high-quality output regardless of the chosen few-step setting. Since adversarial distillation objectives (Sauer et al., 2024b; Lin et al., 2024) are known to preserve sharpness and fine details better than DMD under low-step constraints, this alignment is performed using a discriminator $D_{\psi}^{n}$. This discriminator’s task is to make the outputs of adjacent denoising steps indistinguishable through a relativistic pairing GAN objective (Jolicoeur-Martineau, 2018). The adversarial self-distillation loss is represented as:
@@ 3.3 First-Frame Enhancement
Existing causal video generation methods typically employ a uniform number of denoising steps for each frame. However, in causal generation process, the quality of subsequent frames is critically dependent on the preceding frames. The quality defects in earlier frames are highly likely to be propagated throughout the rest of the sequence. This is particularly true for the first frame, which has no prior context and must synthesize a high-quality initial state from a zero-data starting point. Consequently, the generation of the first frame requires dedicated attention to mitigate accumulated error and ensure high overall video quality. Fig. 4 shows the similarity of predicted $\hat{x}_{0}$ for different video frames across various denoising steps. It reveals that the first frame has low similarity across steps, indicating that each denoising step is crucial. In contrast, subsequent frames exhibit higher similarity, suggesting a greater redundancy that makes them more suitable for a few-step prediction. These results are based on the Self Forcing model and are consistent across multiple prompt variants and random seeds. We include additional results across scenarios in the Appendix Appendix˜C.
Based on this observation, we propose a novel First-Frame Enhancement (FFE) denoising strategy, as detailed in Alg. 2. The first frame undergoes a more intensive denoising process, requiring a minimum of four steps, while subsequent frames can be generated with a significantly reduced number of steps, such as one or two. This frame-based control allows us to enhance the quality of the generated video in the few-step scenario.
@@ 4.1 Implementations
Our approach employs a causal video distillation framework based on the Self Forcing (Huang et al., 2025) training paradigm. The causal model architecture builds upon the Wan2.1-T2V-1.3B (Wang et al., 2025) backbone, a Flow Matching based model (Lipman et al., 2023). We adopt CausVid’s (Yin et al., 2025) initialization protocol to stabilize early causal training phases via asymmetric distillation from a pre-trained bidirectional teacher model. For training data, we utilize the exact text prompts from Self Forcing for fair comparison. The diffusion process is optimized with a 4-step denoising schedule during training, while using our step-skipping strategy during inference to accelerate generation. Adversarial training objective is incorporated via integration of RpGAN (Jolicoeur-Martineau, 2019) objectives with R1 (Mescheder et al., 2018) and R2 regularization terms following R3GAN (Huang et al., 2024a). And we use the frozen fake score function as the backbone of the discriminator following DMD2 Yin et al. (2024a). To reduce computational cost and the number of parameters, we model discriminators $D^{n}$’s output logits as the n-th dimensional output logit of the final layer, thus different $D^{n}$ share the same backbone and classifier head. The performance is rigorously assessed using VBench (Huang et al., 2024b) for multidimensional evaluation and user preference studies to quantify human-perceived visual quality and semantic alignment.
@@ 4.2 Comparison with Existing Baselines
Table 1: Comparison with relevant baselines. We compare our method with representative open-source video generation models of similar parameter sizes and resolutions. The n-step generation with FFE strategy is denoted with $n^{*}$. denotes the retrained version under 1 & 2 step settings.
@@
Figure 6: User preference study. “(n)” == number of denoising steps. denotes the retrained version under 1 & 2 step settings.
@@
As illustrated in Fig. 5, our method is capable of generating high-fidelity videos using only approximately 8% and 13% of the denoising steps required by Wan 2.1(Wang et al., 2025) and SkyReels(Chen et al., 2025), respectively, while achieving visually superior results. This leads to a substantial reduction in computational cost. Moreover, under the same number of denoising steps, our approach yields improved detail and visual quality compared to CausVid(Yin et al., 2025) and better video details compared to Self-forcing(Huang et al., 2025). Fig. 6 shows the user study results comparing our model against several important baselines. Our approach is consistently preferred over many-step baselines, including the many-step Wan2.1 that our model is initialized from. Meanwhile, our results are 70% better than CausVid (Yin et al., 2025) and on par with Self Forcing, since both our model and Self Forcing use identical DMD-based supervision in the 4-step setting. For the extreme 1-step and 2-step generation setting, our method yields a significantly 96% and 62% preference over Self-Forcing, respectively. This demonstrates the effectiveness of our proposed ASD and FFE strategy in improving extremely few-step video generation.
@@ 4.3 Ablation Study
To better understand the contribution of each component to skip-step generation, we conducted a comprehensive ablation study, as shown in Tab. 2. As shown in the first two rows, using identical training data, our ASD training method consistently improves the quality of both 1-step and 2-step generation during inference. For example, it raises the Total Score by 2.52 and the Semantic Score by 6.84 under one-step generation. Comparison between the first and third rows reveals that the FFE inference strategy leads to substantial gains in overall video quality—notably, a 4.19 increase in Total Score and a 10.64 improvement in Semantic Score for one-step generation, even exceeding the quality of the original two-step generation results (row 1). Moreover, combining both methodologies further enhances performance, as demonstrated in the last row.
@@ 7 Ethics Statement
This work focuses on improving the efficiency of diffusion-based video generation through distillation. Our contributions are methodological and technical in nature, and do not involve human subjects, personal or sensitive data, or medical applications. The experiments are conducted on publicly available benchmark datasets (e.g., VBench), which are widely used in prior research and do not contain personally identifiable information.
We confirm that this research adheres to the ICLR Code of Ethics and complies with all relevant institutional and legal standards. No conflicts of interest or external sponsorship influencing the work are present.
8 Acknowledgements
This work was supported by the National Natural Science Foundation of China under grant 62372341. The authors were partially supported by the Intelligent Computing Center of the National Cybersecurity Talent and Innovation Base, Wuhan.
@@ Appendix A Details of Implementations
where $x^{n+1}_{t}\sim p_{\theta,\text{n+1},t}$, $x^{n}_{t}\sim p_{\theta,\text{n},t}$ are the noisy (n+1)-step generation data and n-step generation data, respectively, $\epsilon$ and $\hat{\epsilon}$ are Gaussian noise sampled from $\mathcal{N}(0,1)$, and $f_{\psi}$ is the critic network (discriminator) of GAN. We use $\lambda=600$, $\sigma=0.05$ for all experiments. The generator, fake score estimator, and discriminator are optimized alternately with a ratio of 1:4:1, following(Yin et al., 2024a; Huang et al., 2025). Following Self-Forcing, we inserted additional cross-attention layers and classification heads (serving as discriminator heads) at layers 12, 21, and 29 of the fake score model. These discriminator heads operate solely on backbone features and query tokens, without access to the timestep t. This implementation detail will be explicitly included in the revised manuscript.
Note that we discard ASD Loss for the last step denoising. For simplicity, we reserve the index n of the n-th discriminator $D_{n}$ that discriminates the $n$-step generation distribution from that of $(n+1)$-step here.
For the retrained Self-Forcing baselines, we use exactly the same hyperparameters as the official implementation, except for changing the student model’s number of denoising step list from [1000, 750, 500, 250] to [1000, 500] or [1000]. Specifically, we used: real score CFG weight: 3.0, optimizers: AdamW for both generator and discriminator with $\beta_{1}$ = 0, $\beta_{2}$ = 0.999, $\epsilon$ = 1e-8, weight decay = 0.01, learning rate (generator): 2e62e-6, learning rate (discriminator): 4e74e-7, generator/discriminator update ratio: 5:1. Training is monitored until convergence.
Appendix B VBench Scores Across All Dimensions
Figure 8: VBench scores visualization. We compare our results with Self Forcing Huang et al. (2025) variants using all 16 VBench metrics under 1-/2-step setting. The retrained 1-/2-step version denotes as Self Forcing.
In Figure˜8, we evaluate our method and Self-Forcing across all 16 VBench metrics under both 1-step and 2-step generation settings. Our method consistently outperforms Self Forcing, achieving notably higher scores in semantic alignment-particularly in object class, multiple objects, spatial relationships, and scene. Moreover, it demonstrates superior performance in dynamic modeling, as evidenced by a significantly higher dynamic degree score.
Appendix C Causal Denoising Trajectory Similarity Across Scenarios
Figure 9: Cosine similarity matrices of predicted $\hat{x_{0}}$ across denoising steps (1 to 4) in causal diffusion video generation, comparing our method and Self-Forcing (Huang et al., 2025). We report results averaged separately for VBench, complex object motion, and high camera motion.
To further validate the generality of the phenomenon in Section˜3.3, we apply the analysis to both our method and Self-Forcing across three diverse scenarios: VBench(Huang et al., 2024b), complex object motion, and high camera motion. For each, we record denoising trajectories from 200 video samples, yielding Figure˜9. The figure shows that, across all scenarios, the first frame exhibits low similarity across denoising steps, indicating each step is critical. In contrast, subsequent frames show high inter-step similarity, suggesting greater redundancy and thus greater suitability for few-step prediction. Moreover, our method achieves even higher similarity in later frames, demonstrating that ASD enables the student to match multi-step quality with fewer steps, effectively supporting step-skipping inference.
Appendix D Reduced Distributional Gaps Between Adjacent Denoising Steps
To empirically validate the smoother transitions between adjacent denoising steps, we conduct a quantitative analysis using the Fréchet Video Distance (FVD). Table 3 presents FVD values computed between outputs at step $n$ and step $n+1$ (i.e., adjacent steps), as well as between step $n$ and the 50-step 14B teacher model.
Table 3: Fréchet Video Distance (FVD) between video distributions at different denoising steps. Lower FVD values indicate smaller distributional gaps.
Comparison Step 1 Step 2 Step 3 Step 4
$n\to n+1$ 732 1136 441 N/A
$n\to\text{teacher}$ 1836 1646 1454 1448
As illustrated in Table 3, the FVD between adjacent denoising steps is consistently lower than that between each step and the final teacher output. For instance, at step 1, the adjacent-step FVD is 732, much smaller than 1836 for the teacher-step comparison. These results provide quantitative support that the distributional discrepancy between consecutive steps is substantially smaller than the divergence from the teacher distribution. The Adjacent-Step Distribution (ASD) objective capitalizes on this property by aligning each step with its immediate successor, which exhibits lower distributional divergence. Overall, these findings corroborate the core assumption of our approach: transitions between adjacent denoising steps are indeed smoother.
Appendix E Few-step Long-video Generation Evaluation
To demonstrate that our method enables high-quality, long-video generation with few (1 or 2) denoising steps—and effectively mitigates error accumulation—we generate 20-second continuations using both 1-step and 2-step inference for our approach and Self-Forcing, and evaluate them on VBench(Huang et al., 2024b).
As shown in Table˜4, our method overall outperforms Self-Forcing both quantitatively and qualitatively in long-video generation. For visual comparison, Figure˜10 presents side-by-side 2*-step (ours) vs. 2-step (Self-Forcing) 20-second samples. Our results exhibit clearly superior performance in motion dynamics (rows 3, 4, 6), visual detail and fidelity (rows 1, 2, 3, 6), and camera motion realism (rows 1, 2, 4).
These results confirm that ASD not only reduces computational cost but also enhances temporal coherence and visual quality in extended video generation.
Table 4: VBench evaluation scores for our method vs. Self-Forcing under few-step long-video (20-second) inference settings
Method Total Score Quality Score Semantic Score
Self Forcing(Huang et al., 2025) (2-step) 0.8250 0.8293 0.8076
Ours (2*-step) 0.8263 0.8329 0.7998
Self Forcing(Huang et al., 2025) (1-step) 0.8066 0.8101 0.7923
Ours (1*-step) 0.8200 0.8248 0.8011
Figure 10: Our method versus Self-Forcing on uniformly sampled frames (including first and last) from 20-second long videos.
Appendix F Hyperparameter Analysis